# EVM-QuestBench: An Execution-Grounded Benchmark for Natural-Language Transaction Code Generation

### 10 Mar 2026 v2

@@ Abstract
Large language models are increasingly applied to various development scenarios. However, in on-chain transaction scenarios, even a minor error can cause irreversible loss for users. Existing evaluations often overlook execution accuracy and safety. We introduce EVM-QuestBench, an execution-grounded benchmark for natural-language transaction-script generation on EVM-compatible chains. The benchmark employs dynamic evaluation: instructions are sampled from template pools, numeric parameters are drawn from predefined intervals, and validators verify outcomes against these instantiated values. EVM-QuestBench contains 107 tasks (62 atomic, 45 composite). Its modular architecture enables rapid task development. The runner executes scripts on a forked EVM chain with snapshot isolation; composite tasks apply step-efficiency decay. We evaluate 20 models with 5 independent rounds each and find large performance gaps, with split scores revealing persistent asymmetry between single-action precision and multi-step workflow completion. Code: https://anonymous.4open.science/r/bsc_quest_bench-A9CF/.
@@ 1 Introduction
We study natural language transaction script generation on EVM-compatible chains. The target output is a client-side TypeScript module that composes calls to deployed contracts. The module returns unsigned transaction payloads. A standardized runner signs and executes the requests on a forked chain. Validators check receipts and post-state constraints. This design measures end-to-end correctness under execution.
@@
  • We release EVM-QuestBench, a benchmark for natural language to transaction script generation on EVM-compatible chains, with Atomic and Composite splits.
@@
  • We report results on 20 models with 5-round statistical analysis, including confidence intervals and rank-order consistency, that separates single-action precision from multi-step workflow completion.
@@ 2 Related Work
A major limitation of reference-based code evaluation is that surface overlap can reward outputs that resemble references but fail functionally. Execution-based benchmarks address this by running generated programs against tests, including HumanEval and MBPP for function synthesis (chen2021evaluating; austin2021program), and extensions to library usage and multilingual settings (wang2023odex; khan2024xcodeeval; yan2024codescope). However, most prior benchmarks focus on stateless, sandboxed functions and do not model shared external state or irreversible actions.
Real-world software engineering and multi-step interaction.
SWE-bench evaluates issue resolution by applying patches to real repositories and validating with test suites (jimenez2024swebench), emphasizing repository-scale context and cross-file edits. Separately, agent-style benchmarks study multi-step tool use and action sequencing in interactive environments (qin2023toollm; liu2023agentbench). These lines are complementary, but they typically do not target blockchain-specific constraints such as transaction construction, revert conditions, strict unit handling, or protocol-imposed prerequisites.
Blockchain-specific evaluation and our positioning.
Blockchain interaction introduces shared mutable state, protocol prerequisites, and revert risk that are absent in standard code tasks. Prior blockchain benchmarks (e.g., Solana-focused transaction evaluations) (solana2025bench) do not explicitly disentangle single-transaction precision from multi-transaction workflow completion under a unified execution and validation interface. EVM-QuestBench fills this gap by evaluating natural-language-to-transaction-script generation on EVM-compatible chains with atomic/composite splits. Tasks are specified declaratively as JSON with reusable validator components, enabling new tasks to be added with minimal engineering effort while preserving comparable execution-grounded evaluation.
@@ 3.1 Benchmark Overview
Figure 1: EVM-QuestBench evaluation architecture. Tasks are defined as JSON specs, dynamically instantiated into concrete instances, executed through a fixed TypeScript interface, and scored by validator-driven post-state checks. Composite tasks additionally apply step-efficiency decay.
@@ 3.2 Task Specification and Data Format
A typical instance provides an instruction such as “Swap 0.1 ETH to USDT” together with an execution context (RPC endpoint, agent address, and a contract address map). The model returns a module whose executeSkill function emits a transaction request that calls the router with a concrete to address and data calldata. The validator then checks transaction success and verifies the expected balance change within the task tolerance (Section 3).
@@ 3.3 Benchmark Composition
EVM-QuestBench contains 107 tasks (62 atomic, 45 composite) with a maximum total score of 10,700. Figure 2 shows the split sizes and Figure 3 shows the composite workflow complexity.
@@ 3.4 Difficulty and Coverage
Each task is annotated with a difficulty label from {easy, medium, hard}. Atomic tasks are dominated by medium (56.5%), while composite tasks skew harder (hard: 24.4%), reflecting additional constraints from prerequisite handling and step ordering. Atomic tasks span subcategories including ERC20 operations, native transfers, NFT operations, swaps, staking, and queries. Composite tasks diversify by workflow motifs: batch operations, swap-centric workflows, liquidity workflows, staking workflows, and query-and-verify patterns.
@@ 3.5 Instruction Templates and Parameterization
Each task provides multiple natural language templates that render a user goal as an instruction, together with a parameter schema that binds slots in the template (e.g., token symbols, contract keys, recipient addresses, and amounts). These templates are generated by an LLM during benchmark development to ensure diverse and realistic phrasing. Atomic tasks include 3355 templates (mean $3.97$, median $4$), while composite tasks include 2244 templates (mean $2.82$, median $3$).
At evaluation time, the runner randomly selects one template from the candidate pool and samples numeric parameter values uniformly within predefined ranges specified in the task configuration. For both atomic and composite tasks, all numeric values (e.g., transfer amounts, token quantities, percentages) are dynamically generated within configured intervals rather than fixed. The runner injects these sampled values into the selected template and passes the resolved parameters to validators.
Benefits of dynamic parameterization.
Dynamic numeric sampling provides several evaluation advantages over fixed test cases. First, it prevents trivial memorization: models cannot succeed by caching specific amount patterns from training data. Second, it tests numeric reasoning robustness: models must correctly parse arbitrary values from natural language (e.g., “transfer 0.37 tokens” vs. “transfer 100 tokens”) and propagate them through unit conversions and calldata encoding. Third, it exposes edge-case failures: boundary values and unusual magnitudes reveal precision errors, overflow handling issues, and decimal truncation bugs that fixed test cases might miss. Fourth, it enables statistical significance: multiple runs with different parameter samples yield confidence intervals rather than single-point estimates. This setup ensures that models must generalize across different numeric inputs rather than pattern-match against known examples. Correctness depends on consistent use of task-provided context, including the contract address map, token decimals, and protocol constraints.
@@ 3.6 Validators and Post Execution Constraints
EVM-QuestBench uses validator-based scoring rather than reference code matching. This design enables evaluation of functional correctness under dynamic parameterization: validators receive the same randomly sampled parameters that were injected into the natural language instruction, and verify that the model’s output produces the expected on-chain effects for those specific values.
Validator architecture.
Each atomic task type has a dedicated validator class (e.g., ERC20TransferValidator, SwapExactETHForTokensValidator). At instantiation, validators receive the dynamically sampled parameters from the runner, including target addresses, token amounts, and decimal configurations. Composite tasks use a CompositeValidator that loads and delegates to the appropriate atomic validator based on the key operation in the workflow.
Validation against dynamic parameters.
The validator converts human-readable parameters to chain-native units (e.g., $0.1$ tokens $\rightarrow$ $0.1\times 10^{18}$ wei for an 18-decimal token) and compares against actual on-chain state changes. For a transfer task, the validator checks: (i) the sender’s balance decreased by the expected amount, and (ii) the receiver’s balance increased by the same amount. This approach ensures that correctness is measured against the specific numeric values presented in the instruction, not against a fixed reference output.
Weighted scoring components.
Validators implement protocol and task-specific checks over receipts and post-state. A typical atomic validator (e.g., ERC20 transfer) computes a weighted score from four check categories: transaction success (30 points), contract address correctness (20 points), function signature correctness (20 points), and state change verification (30 points). Each check is binary pass/fail with its assigned weight. The final score is the sum of passed check weights.
Tolerance handling.
Many tasks allow small tolerances to account for AMM slippage or rounding effects. Transfer validators typically allow a $1$ to 22% relative tolerance on amounts. Swap and liquidity operations use a slippage-style tolerance (commonly 55%). Approval tasks require an exact allowance match. These tolerances are applied during the state change verification step, ensuring that minor deviations from expected values (due to protocol mechanics) do not penalize otherwise correct solutions.
Composite validation.
Composite validators compute a base success score from the final end-state condition, then apply step efficiency decay as described in Section 4. The validator loads the composite task definition, identifies the key operation, and reuses the corresponding atomic validator to score the final transaction. This modular design eliminates code duplication and ensures scoring consistency between atomic and composite evaluations of the same operation type.
@@ 4.2 Execution Environment
Evaluation runs on an Anvil fork of BSC mainnet (chain ID 56), an EVM-compatible chain. We use BSC as the instantiation; the benchmark design generalizes to other EVM chains. The runner specifies the upstream RPC endpoint. The fork block height is not pinned; each evaluation uses the latest block at fork time. Within an evaluation run, we enforce comparability across tasks via snapshot isolation (Section 4.3).
@@ 5.1 Leaderboard
We evaluate 20 models on EVM-QuestBench with 5 independent evaluation rounds per model (totaling $5\times 107\times 20=10{,}700$ task executions). Each round uses freshly sampled numeric parameters, independent LLM calls (temperature =0.7=0.7), and clean snapshot-isolated environments. Table 1 reports the mean Atomic, Composite, and Total scores across 5 rounds, along with standard deviation (SD), 95% confidence interval (CI, computed as mean $\pm\,t_{0.025,4}\times\text{SE}$ with $t=2.776$ for df =4=4), and coefficient of variation (CV%).
The top three models exceed 7,7007{,}700 mean total points. Claude-Sonnet-4.5 achieves the highest mean total (8,2368{,}236) with low variance (CV $2.1\%$). 15 out of 20 models exhibit CV% below $9\%$, demonstrating that evaluation scores are robust to the combined variance from dynamic parameter sampling, temperature-based generation, and RPC execution.
# Model Atom Comp Total SD CV% 95% CI
1 Claude-Sonnet-4.5 4278 3958 8236 174 2.1 [8020, 8451]
2 Gemini-3-Pro 4302 3561 7863 267 3.4 [7531, 8195]
3 GPT-5 4128 3646 7774 407 5.2 [7269, 8279]
4 GPT-5.1 3715 3617 7332 144 2.0 [7154, 7510]
5 Kimi-K2-Thinking 3605 3633 7238 207 2.9 [6982, 7495]
6 Gemini-2.5-Flash 3265 3768 7033 127 1.8 [6876, 7191]
7 GPT-OSS-120B 3458 3468 6926 436 6.3 [6384, 7467]
8 DeepSeek-V3.2 3071 3638 6709 384 5.7 [6232, 7186]
9 Grok-4-Fast 2980 3207 6187 224 3.6 [5909, 6465]
10 Qwen3-235B 3609 2116 5724 630 11.0 [4943, 6506]
11 GPT-5.1-Mini 2510 3190 5700 510 8.9 [5067, 6333]
12 Qwen3-Max 2775 2883 5658 229 4.1 [5374, 5943]
13 Claude-Haiku-4.5 3221 2367 5588 316 5.7 [5196, 5981]
14 Mimo-V2-Flash 3189 1723 4912 218 4.4 [4641, 5183]
15 GLM-4.6 2881 1971 4852 295 6.1 [4486, 5218]
16 Qwen3-30B 1554 2011 3565 398 11.2 [3072, 4059]
17 Devstral-2512 2892 21 2914 303 10.4 [2537, 3290]
18 Qwen3-Coder 1801 160 1961 291 14.8 [1600, 2322]
19 Qwen3-Coder-Flash 223 38 260 212 81.4 [3-3, 524]
20 Qwen3-Coder-30B 140 5 145 61 42.0 [69, 221]
Table 1: Leaderboard (avg@5). All scores are means over 5 independent rounds. Atom = Atomic (62 tasks, max 6,200), Comp = Composite (45 tasks, max 4,500), Total = 107 tasks (max 10,700). Spearman’s rank correlation between any two rounds averages $\rho=0.960$ (all $p<0.001$, range: 0.940–0.983).
@@ 5.2 Atomic versus Composite Capability Separation
Figure 4 visualizes the Atomic–Composite plane (a full ranking bar chart is in Appendix E, Figure 6). Models spread widely, indicating that single-step correctness and multi-step completion are only partially coupled.
@@ 5.3 Composite Execution Quality
Table 2 reports fine-grained composite metrics averaged over 5 rounds. We analyze step efficiency ($\bar{K}_{\text{act}}$, Eff%) as primary dimensions.
Model Score Pass $\bar{K}$ Eff%
Claude-Sonnet-4.5 3958 41.0 3.8 88.0
Gemini-2.5-Flash 3768 38.6 4.1 83.7
GPT-5 3646 38.0 3.6 83.8
DeepSeek-V3.2 3638 38.2 3.5 84.5
Kimi-K2-Thinking 3633 33.6 3.6 84.2
GPT-5.1 3617 38.0 4.1 80.4
Gemini-3-Pro 3561 37.6 4.3 79.7
GPT-OSS-120B 3468 35.2 3.9 82.6
Grok-4-Fast 3207 32.0 4.4 76.4
GPT-5.1-Mini 3190 31.0 4.7 73.6
Qwen3-Max 2883 25.8 4.9 71.5
Claude-Haiku-4.5 2367 13.0 5.9 60.7
Qwen3-235B 2116 11.0 5.8 62.6
Qwen3-30B 2011 8.4 6.1 58.8
GLM-4.6 1971 10.2 6.0 58.3
Mimo-V2-Flash 1723 9.0 6.2 56.9
Qwen3-Coder 160 0.4 6.7 53.3
Qwen3-Coder-Flash 38 0.2 6.8 51.1
Devstral-2512 21 0.0 6.8 50.8
Qwen3-Coder-30B 5 0.0 6.8 50.6
Table 2: Composite execution quality (avg@5, 45 tasks). Score and Pass are means over 5 rounds. $\bar{K}$ = mean $K_{\text{act}}$, Eff = step efficiency.
@@
5.4 Where Workflows Fail
Figure 5 reports pass rates by workflow pattern. The hardest patterns are multi stage DeFi workflows combining liquidity and staking, as well as batch approval patterns. These concentrate three failure surfaces: (i) prerequisite correctness, (ii) cross step parameter consistency, and (iii) execution robustness under multiple transactions. Additional analysis is in Appendix E.

### 16 Mar 2026 v3

@@ Abstract
Large language models are increasingly applied to various development scenarios. However, in on-chain transaction scenarios, even a minor error can cause irreversible loss for users. Existing evaluations often overlook execution accuracy and safety. We introduce EVM-QuestBench, an execution-grounded benchmark for natural-language transaction-script generation on EVM-compatible chains. The benchmark employs dynamic evaluation: instructions are sampled from template pools, numeric parameters are drawn from predefined intervals, and validators verify outcomes against these instantiated values. EVM-QuestBench contains 107 tasks (62 atomic, 45 composite). Its modular architecture enables rapid task development. The runner executes scripts on a forked EVM chain with snapshot isolation; composite tasks apply step-efficiency decay. We evaluate 20 models with 5 independent rounds each and find large performance gaps, with split scores revealing persistent asymmetry between single-action precision and multi-step workflow completion.
@@ 1 Introduction
  • We release EVM-QuestBench, a benchmark for natural language to transaction script generation on EVM-compatible chains, with Atomic and Composite splits. Its declarative task design keeps evaluation costs low: each task round consumes a small number of tokens, enabling comprehensive model assessment at minimal token cost.
@@
  • We report results on 20 models with 5-round statistical analysis, including standard deviations, observed ranges, and rank-order consistency, that separates single-action precision from multi-step workflow completion.
@@ 2 Related Work
Reference-based code evaluation metrics such as BLEU and CodeBLEU (papineni2002bleu; ren2020codebleu) share a fundamental flaw: they measure surface similarity rather than functional correctness, so code that resembles a reference but fails at runtime can still receive a high score. Execution-based benchmarks address this by running generated programs against tests, including HumanEval and MBPP for function synthesis (chen2021evaluating; austin2021program), and extensions to library usage and multilingual settings (wang2023odex; khan2024xcodeeval; yan2024codescope). However, most prior benchmarks focus on stateless, sandboxed functions and do not model shared external state or irreversible actions.
Real-world software engineering, agent benchmarks, and blockchain constraints.
SWE-bench evaluates issue resolution by applying patches to real repositories (jimenez2024swebench), and agent-style benchmarks study multi-step tool use in interactive environments (qin2023toollm; liu2023agentbench). However, these benchmarks do not target the constraints that characterize blockchain interaction: shared mutable state, protocol prerequisites, strict unit and decimal handling, and irreversible revert risk. Prior blockchain benchmarks such as Solana-focused transaction evaluations (solana2025bench) also do not disentangle single-transaction precision from multi-transaction workflow completion under a unified execution and validation interface. EVM-QuestBench fills this gap by evaluating natural-language-to-transaction-script generation on EVM-compatible chains with atomic/composite splits, using declaratively specified tasks and reusable validator components that keep development costs low while preserving execution-grounded evaluation.
@@ 3.1 Benchmark Overview
Figure 1: EVM-QuestBench evaluation architecture and end-to-end pipeline. A natural language instruction is sampled from a template pool with dynamic numeric parameters, passed to the LLM for TypeScript script generation, executed on a snapshot-isolated forked chain, and scored by task-specific validators against post-state constraints. Composite tasks additionally apply step-efficiency decay.
@@ 3.2 Task Specification and Data Format
A typical instance provides an instruction such as “Swap 0.1 BNB to USDT” together with an execution context (RPC endpoint, agent address, and a contract address map). The model returns a module whose executeSkill function emits a transaction request that calls the router with a concrete to address and data calldata. The validator then checks transaction success and verifies the expected balance change within the task tolerance (Section 3).
@@ 3.3 Benchmark Composition
EVM-QuestBench contains 107 tasks (62 atomic, 45 composite) with a maximum total score of 10,700. Figure 4 (Appendix B) shows the split sizes and Figure 5 (Appendix B) shows the composite workflow complexity.
@@ 3.4 Difficulty and Coverage
Each task is annotated with a difficulty label from {easy, easy-medium, medium, hard}. Atomic tasks are dominated by medium (56.5%), while composite tasks skew harder (hard: 24.4%), reflecting additional constraints from prerequisite handling and step ordering. Atomic tasks span subcategories including ERC20 operations, native transfers, NFT operations, swaps, staking, and queries. Composite tasks diversify by workflow motifs: batch operations, swap-centric workflows, liquidity workflows, staking workflows, and query-and-verify patterns.
@@ 3.5 Instruction Templates and Parameterization
Each task provides multiple natural language templates (atomic: 3355, mean $3.97$; composite: 2244, mean $2.82$), generated by an LLM to ensure diverse and realistic phrasing. At evaluation time, the runner randomly selects one template and samples numeric parameter values uniformly within predefined ranges, injecting them into the selected template and passing the resolved parameters to validators. Dynamic sampling prevents trivial memorization, forces models to generalize across arbitrary values and unit conversions, and keeps per-run token consumption low by reusing the same 107 tasks with fresh parameters across rounds.
To control for wording bias, all 373 templates were rated on a 1–5 clarity scale by three SOTA models and averaged: 336 are precise (90.1%) and none are vague, confirming the pool is consistently unambiguous. Full details and the difficulty selection API are in Appendix F.
@@ 3.6 Validators and Post Execution Constraints
EVM-QuestBench uses validator-based scoring rather than reference code matching. Validators receive the same dynamically sampled parameters injected into the natural language instruction and verify that the model’s output produces the expected on-chain effects for those specific values.
Each atomic task type has a dedicated validator that converts human-readable parameters to chain-native units (e.g., $0.1$ tokens $\rightarrow$ $0.1\times 10^{18}$ wei) and compares against actual on-chain state changes. A typical atomic validator computes a weighted score from four binary checks: transaction success (30 pts), contract address correctness (20 pts), function signature correctness (20 pts), and state change verification (30 pts). Tolerance thresholds are calibrated per task family: transfer validators apply a 0.10.1% relative tolerance, approval validators require an exact match, and swap/liquidity operations use a 55% slippage tolerance to account for AMM non-determinism. Composite validators score the final end-state condition and apply step efficiency decay (Section 4), reusing the corresponding atomic validator for consistency.
@@ 4 Evaluation Setup
Figure 2: End-to-end evaluation pipeline of EVM-QuestBench. A natural language instruction is instantiated from a template with dynamically sampled parameters, fed to the LLM for TypeScript script generation, executed by the runner on a snapshot-isolated forked chain, and finally scored by task-specific validators against post-execution on-chain state.
@@ 4.2 Execution Environment
Evaluation runs on an Anvil fork of BSC mainnet (chain ID 56), an EVM-compatible chain. We use BSC as the instantiation; the benchmark design generalizes to other EVM chains. The runner specifies the upstream RPC endpoint. The fork block height is not pinned; each evaluation run forks at the latest block. Since all 107 tasks within a run share the same fork origin and snapshot isolation (Section 4.3) prevents cross-task state leakage, intra-run comparability is fully preserved.
@@
Client library backend.
All experiments use ethers.js v6, the de-facto standard for EVM client scripting, chosen for its broad representation in LLM training corpora and as the most equitable cross-model baseline. A viem v2 backend is also supported; switching requires only a different system configuration file with no changes to validators or scoring. Full backend details are in Appendix C.
@@ 5.1 Leaderboard
We evaluate 20 models on EVM-QuestBench with 5 independent evaluation rounds per model (totaling $5\times 107\times 20=10{,}700$ task executions). Each round uses freshly sampled numeric parameters, independent LLM calls (temperature =0.7=0.7), and clean snapshot-isolated environments. Table 1 reports the mean Atomic, Composite, and Total scores across 5 rounds, along with standard deviation (SD), observed range (min, max) over 5 rounds, and coefficient of variation (CV%).
The top three models exceed 7,7007{,}700 mean total points. Claude-Sonnet-4.5 achieves the highest mean total (8,2368{,}236) with low variance (CV $2.1\%$). 14 out of 20 models exhibit CV% below $9\%$, demonstrating that evaluation scores are robust to the combined variance from dynamic parameter sampling, temperature-based generation, and RPC execution.
# Model Atom Comp Total SD CV% Range [min, max] $_{\text{a}} Pass$$_{\text{c}}$ $\bar{K}$ Eff%
1 Claude-Sonnet-4.5 4278 3958 8236 174 2.1 [8017, 8403] 35.0 41.0 3.8 88.0
2 Gemini-3-Pro 4302 3561 7863 267 3.4 [7487, 8195] 36.6 37.6 4.3 79.7
3 GPT-5 4128 3646 7774 407 5.2 [7241, 8121] 31.6 38.0 3.6 83.8
4 GPT-5.1 3715 3617 7332 144 2.0 [7199, 7542] 28.0 38.0 4.1 80.4
5 Kimi-K2-Thinking 3605 3633 7238 207 2.9 [6928, 7498] 29.2 33.6 3.6 84.2
6 Gemini-2.5-Flash 3265 3768 7033 127 1.8 [6837, 7176] 22.0 38.6 4.1 83.7
7 GPT-OSS-120B 3458 3468 6926 436 6.3 [6430, 7633] 26.0 35.2 3.9 82.6
8 DeepSeek-V3.2 3071 3638 6709 384 5.7 [6074, 7088] 21.6 38.2 3.5 84.5
9 Grok-4-Fast 2980 3207 6187 224 3.6 [5913, 6384] 20.8 32.0 4.4 76.4
10 Qwen3-235B 3609 2116 5724 630 11.0 [4735, 6300] 28.0 11.0 5.8 62.6
11 GPT-5.1-Mini 2510 3190 5700 510 8.9 [5139, 6298] 19.4 31.0 4.7 73.6
12 Qwen3-Max 2775 2883 5658 229 4.1 [5388, 5956] 18.8 25.8 4.9 71.5
13 Claude-Haiku-4.5 3221 2367 5588 316 5.7 [5208, 5899] 24.8 13.0 5.9 60.7
14 Mimo-V2-Flash 3189 1723 4912 218 4.4 [4659, 5160] 25.0 9.0 6.2 56.9
15 GLM-4.6 2881 1971 4852 295 6.1 [4522, 5289] 21.2 10.2 6.0 58.3
16 Qwen3-30B 1554 2011 3565 398 11.2 [3212, 4214] 10.8 8.4 6.1 58.8
17 Devstral-2512 2892 21 2914 303 10.4 [2460, 3267] 20.6 0.0 6.8 50.8
18 Qwen3-Coder 1801 160 1961 291 14.8 [1600, 2265] 13.2 0.4 6.7 53.3
19 Qwen3-Coder-Flash 223 38 260 212 81.4 [40, 538] 1.6 0.2 6.8 51.1
20 Qwen3-Coder-30B 140 5 145 61 42.0 [70, 240] 0.6 0.0 6.8 50.6
[…]
@@ 5.2 Atomic versus Composite Capability Separation
Figure 6 (Appendix B) visualizes the Atomic–Composite plane (a full ranking bar chart is in Appendix H, Figure 8). Models spread widely, indicating that single-step correctness and multi-step completion are only partially coupled.
@@ 5.3 Composite Execution Quality
Table 1 (right block) reports fine-grained composite metrics averaged over 5 rounds. We analyze step efficiency ($\bar{K}_{\text{act}}$, Eff%) as primary dimensions.
@@
5.4 Token Usage and Cost Efficiency
Figure 3 plots total token consumption (single run, 107 tasks) against benchmark score, with colour encoding per-run API cost under OpenRouter cache-read pricing.
Figure 3: Total score (avg@5) versus total token usage (single run, 107 tasks). Colour encodes API cost (USD, with prompt caching). Models in the upper-left quadrant achieve high scores with low token budgets.
A full 107-task run costs as little as \$0.06–\$0.29 for standard models (e.g., GPT-OSS-120B at \$0.06, Gemini-2.5-Flash at \$0.29), consuming under 420K tokens. Thinking-enabled models are substantially more expensive (up to \$14.16 for Gemini-3-Pro) due to chain-of-thought token overhead, yet do not consistently outperform efficient non-thinking models. Detailed per-model cost and caching breakdowns are in Appendix L. Workflow failure analysis by pattern is in Appendix H.